# A framework for fluctuating times and counting observables in stochastic excursions

### v1

@@ IV.2 Counting a set of specific transitions
$\beta$ $=CW_{zz^{\prime}}\langle y_{A}|W_{AB}M|z\rangle\langle z^{\prime}|MW_{BA}|x_{A}\rangle,$ (57)
@@
$a_{ji}$ $=C\sqrt{W_{z_{i}z_{i}^{\prime}}W_{z_{j}z_{j}^{\prime}}}\langle y_{A}|W_{AB}|\tilde{M}|z_{i}\rangle\langle z_{j}^{\prime}|\tilde{M}W_{BA}|x_{A}\rangle,$ (60)
@@
Note that these are all of dimension $r$. The rank of the matrices will be the number of transitions that are being monitored. This holds even if the state space itself is of much higher dimension. $\tilde{M}$ is defined as
@@
$P(\hat{Q}_{1},\ldots,\hat{Q}_{r})=\sum_{\rm all\leavevmode\nobreak\ perms}\mathrm{tr}\Big{\{}P_{1}a(P_{1}b)^{n_{1}}(P_{2}b)^{n_{2}}\ldots(P_{r}b)^{n_{r}}\Big{\}},$ (63)
@@
V Fluctuation theorem
The stochastic entropy production during an excursion was defined in Eq. (17). The probability of observing an entropy production $\sigma$ during an excursion starting from $x_{A}$ and ending in $y_{A}$ is given by Eq. (20):
$P_{x_{A}\to y_{A}}(\sigma)=C_{x_{A}\to y_{A}}\int\frac{d\xi}{2\pi}\langle y_{A}|W_{AB\xi}M_{\xi}W_{BA\xi}|x_{A}\rangle e^{-iq\xi},$ (69)
where
@@ V Fluctuation theorem
For the entropy production, the local detailed balance condition (16) implies that the corresponding tilted matrix in Eq. (18) satisfies the special symmetry
$\mathbb{W}_{\xi}^{\rm T}=\mathbb{W}_{-\xi-i}.$ (71)
Changing variables to $-\xi-i\to\xi$ yields
$P_{x_{A}\to y_{A}}(\sigma)=C_{x_{A}\to y_{A}}e^{\sigma}\int\frac{d\xi}{2\pi}\langle x_{A}|W_{AB\xi}M_{\xi}W_{BA\xi}|y_{A}\rangle e^{iq\xi}.$ (72)
The resulting quantity is proportional to $P_{y_{A}\to x_{A}}(-\sigma)$, except for the normalization constant. We therefore conclude that
$\frac{P_{x_{A}\to y_{A}}(\sigma)}{P_{y_{A}\to x_{A}}(-\sigma)}=e^{\sigma}\frac{C_{x_{A}\to y_{A}}}{C_{y_{A}\to x_{A}}}.$ (73)
The coefficient $C_{x_{A}\to y_{A}}/C_{y_{A}\to x_{A}}$ is the ratio between the probability of having an excursion $x_{A}\to y_{A}$ and having its reverse $y_{A}\to x_{A}$. This appears in Eq. (73) because the left-hand side denotes a conditional probability on both ends.
This result is a consequence of the fluctuation theorem at the stochastic level, since the operation of reversing an excursion is bijective and produces another excursion, and the entropy production is odd under this operation. The fluctuation theorem at the trajectory level, as proved in Ref. seifertEntropyProductionStochastic2005a , relates the probabilities of a trajectory $\gamma$ and its reverse $\overline{\gamma}$ to the entropy production rate $\sigma_{\gamma}$ and change of system entropy between the initial and final states $\Delta s_{\gamma}$:
$\frac{P\big{[}\gamma\big{]}}{P\big{[}\overline{\gamma}\big{]}}=e^{\sigma_{\gamma}+\Delta s_{\gamma}}$ (74)
The joint probability of an excursion that starts in $x_{A}$, ends in $y_{A}$, and has entropy production $\sigma$ is
$P[x_{A}\to y_{A},\sigma]=\sum_{\gamma}P[\gamma]\delta_{\sigma_{\gamma},\sigma}\delta_{\gamma_{\text{i}},x_{A}}\delta_{\gamma_{\text{f}},y_{A}},$ (75)
where we sum over all possible excursions, and use three Kronecker deltas: for selecting the excursions associated with entropy production $\sigma_{\gamma}=\sigma$, for selecting the initial state $\gamma_{\text{i}}=x_{A}$ and the final $\gamma_{\text{f}}=y_{A}$. Plugging in Eq. (74),
$P[x_{A}\to y_{A},\sigma]$ $=\sum_{\gamma}P[\overline{\gamma}]e^{\sigma_{\gamma}+\Delta s_{\gamma}}\delta_{\sigma_{\gamma},\sigma}\delta_{\gamma_{\text{i}},x_{A}}\delta_{\gamma_{\text{f}},y_{A}}$
$=e^{\sigma+\Delta s_{x_{A}\to y_{A}}}\sum_{\gamma}P[\overline{\gamma}]\delta_{\sigma_{\gamma},\sigma}\delta_{\gamma_{\text{i}},x_{A}}\delta_{\gamma_{\text{f}},y_{A}}$
$=e^{\sigma}\frac{p_{x_{A}}^{\rm ss}}{p_{y_{A}}^{\rm ss}}P[y_{A}\to x_{A},-\sigma].$ (76)
Finally, noticing that the joint probability can be expressed as a conditional times the marginal
$P[x_{A}\to y_{A},\sigma]=P_{x_{A}\to y_{A}}(\sigma)P[x_{A}\to y_{A}],$ (77)
and that the normalization constant $C$ has the alternative interpretation
$C_{x_{A}\to y_{A}}^{-1}=P[x_{A}\to y_{A}]\frac{\mathcal{K}_{AB}}{p_{x_{A}}^{\rm ss}},$ (78)
we arrive at the result Eq. (73).
Nonetheless, the devised framework for calculating counting observables of excursions yields a direct verification of the fluctuation theorem specialized to these observables. If the considered excursions start and end in the same state, which is the natural scenario when $|A|=1$, the fluctuation theorem simplifies to
$\frac{P_{x_{A}\to x_{A}}(\sigma)}{P_{x_{A}\to x_{A}}(-\sigma)}=e^{\sigma}.$ (79)
Note how this is an FT of the “exchange” type, in the sense that it involves the same probability in the numerator and denominator. This is not an asymptotic formula, since it holds exactly for individual excursions.
@@ VI.1 Simplified absorption refrigerator
Quantum absorption refrigerators have been described in the literature for many years Mitchison2019 ; Levy2012 ; Levy2012a ; Correa2014 ; Correa2013 ; Cangemi2024 . Recently, they have received a surge of interest due to its experimental implementation aamirThermallyDrivenQuantum2025 ; blokQuantumThermodynamicsQuantum2025 ; Campbell2025 . The system consists of three qubits, each connected to their respective reservoirs which are assumed to be bosonic (at different temperatures), denoted $c$ (cold), $h$ (hot), and $w$ (work). The desired task is to cool down the cold qubit, so the temperatures are taken to obey $T_{c}<T_{h}<T_{w}$. Each qubit has a different energy gap given by $\omega_{\alpha}$ where $\alpha=c,\leavevmode\nobreak\ h,\leavevmode\nobreak\ w$. The three qubits interact resonantly (i.e. when $\omega_{c}+\leavevmode\nobreak\ \omega_{w}=\leavevmode\nobreak\ \omega_{h}$) through an effective three-body interaction, allowing the conversion of one excitation in the work and cold reservoirs into one in the hot reservoir, effectively cooling the cold reservoir in the process, see Fig. 2(a). The behavior of this system at the ensemble level is characterized by the cooling window, which provides bounds to the temperatures of the reservoirs such that the system works as a refrigerator:
@@
Figure 2: (a) Diagram of the three-qubit absorption refrigerator, each qubit is connected to its respective heat reservoir and has a different energy gap. Labels are $c$ for cold, $h$ for hot, and $w$ for work. The resonant interaction $g$ couples the states $|010\rangle\leftrightarrow|101\rangle$ and takes place only if $\omega_{c}+\omega_{w}=\omega_{h}$. (b) Energy landscape of the three-qubit refrigerator model, where the arrows indicate the possible transitions. Colors indicate which reservoir induced the transition. Region $A$ consists of the ground state (gray) and $B$ of all excited states (colored). We adopt the convention “$chw$” to denote the states.
[…]
The transition among states come from two sources, the reservoirs whose transitions rates are described by $\Gamma_{\alpha}n_{\alpha}$ for injections and $\Gamma_{\alpha}(n_{\alpha}+1)$ for extractions, and the three body interaction $g^{\prime}$ that couples the states $10\leftrightarrow 10$ 101 provided that $\omega_{h}=\omega_{c}+\omega_{w}$. Since the transitions are induced by a single reservoir, we omit the index $\ell$ in our discussion. $n_{\alpha}=\left[\exp({\omega_{\alpha}}/T_{\alpha})+1\right]^{-1}$ denotes the occupation number and $\Gamma_{\alpha}$ is the coupling strength between the qubit and the reservoir. The stochastic transition matrix (without the diagonal elements) is then given by
@@
Given the topology of the states and transitions in $B$, this model inherits a nice feature when it comes to constructing the counting observables from Eq. (15). Namely, the thermodynamical currents such as cold, hot and work currents can only have three possible values per excursion. To see this, consider the cold current that is constructed by setting the weights $\nu_{xy}=\pm 1$ in Eq. (15) if the transition $x\rightarrow y$ removes (resp. adds) an excitation from (resp. into) the cold reservoir. The three possibilities for the cold current are: a success where cooling took place and the cold current is $\hat{J}_{c}=1$, e.g. $000\to 100\to 101\to 010\to 000$; a disaster where heating took place the cold current is $\hat{J}_{c}=-1$, e.g. $000\to 010\to 101\to 001\to 000$, and a bounce/fail where $\hat{J}_{c}=0$, e.g. $000\to 100\to 000$. This feature also arises for the entropy production, where the positive and negative contributions are related via fluctuation theorems (73).
@@ VI.2 Full absorption refrigerator
[…]
@@
[…]
$P(\hat{J}_{c})=P_{\textrm{suc}}e^{i\xi}+P_{\textrm{dis}}e^{-i\xi}+P_{\textrm{fail}}$ (83)
@@
where the coefficients are the probabilities that the counting field variable has picked up a particular phase. In other words, $P_{\textrm{suc}}$ is the probability that the cold current has value $\hat{J}_{c}=1$ in an excursion. Likewise, $P_{\textrm{dis}}$ is the probability that the cold current has value $\hat{J}_{c}=-1$ and $P_{\textrm{fail}}$ the probability of $\hat{J}_{c}=0$. Fig. 3(a) shows the behavior of the probabilities as a function of the cold bath occupation number $n_{c}$, where it is clear that cooling failures (bounces) dominate the dynamics. The value where $P_{\rm suc}$ becomes larger than $P_{\rm dis}$ is precisely the value of $n_{c}$ where the cooling window begins. In Fig. 3(b), we show the covariance between the cold current $\hat{J}_{c}$ and the excursion duration $\hat{\mathcal{T}}$ as a function of the cold bath occupation $n_{c}$. Interestingly, longer excursions tend to have larger $\hat{J}_{c}$ in the cooling window, which is not necessarily the case when $\langle\hat{J}_{c}\rangle>0$, and once again, a vanishing covariance pinpoints the value of $n_{c}$ where the cooling window begins.
@@
Equation (83) illustrates another feat of the excursion framework. This is a new paradigm for analyzing the performance of cooling/heating protocols beyond the steady state. The steady state current would take into account only the difference between the probability of successes and disasters, overlooking how dominant failures are. This raises questions for designing useful thermodynamic protocols, e.g. a protocol may have a large cooling capacity $P_{\textrm{suc}}\gg P_{\textrm{dis}}$ but have low cooling precision $P_{\textrm{fail}}\gg P_{\textrm{suc}}$. Non-trivial tradeoffs emerge from this analysis.
@@
For the next two counting observables we discuss, we go back to the case where region $A$ is spanned only by the state $000$. This allow us to use the results from Eqs. (44) and (45) to characterize fluctuations. Entropy production $\hat{\Sigma}$ was constructed in Eq. (17), and we highlighted its importance in our framework when we discussed fluctuation theorems in Sec. V. Likewise, the dynamical activity $\hat{\mathcal{A}}$ was constructed in Eq. (46) and we have provided analytical formulas for its expectation value in Eq. (53) and variance in Eq. (IV.1). In Figs. 4(a),(b) we characterize the diffusion coefficient (45) in terms of its components, both for the entropy production $D_{\Sigma}$ and the dynamical activity $D_{\mathcal{A}}$. In the former, the fluctuations are strongly dominated by the variance term of $\hat{Q}=\hat{\Sigma}$; whereas in the latter there is a clear competition between different contributions, with $E(\hat{\mathcal{A}})^{2}$ being the most relevant term. This result sheds light on the very nature of noise in FCS, where the dominant contribution can change depending on the counting observable. $\hat{\Sigma}$ is an observable that can have positive and negative values related by fluctuation theorems, whereas $\hat{\mathcal{A}}$ is strictly positive. The stark contrast of behavior on the two observables shows the versatility of the excursion framework to characterize fluctuations.
@@
Figure 4: Diffusion coefficient and its decomposition for (a) entropy production, and (b) dynamical activity as a function of $g$, with $n_{c}=0.1$ and $\Gamma_{c}=0.1$. For both plots, we fixed $\Gamma_{h}=50,\ \Gamma_{w}=0.5,\ n_{h}=0.005,\ n_{w}=0.5$. Parameters are well in accordance with the parameter regime of aamirThermallyDrivenQuantum2025 and the classical approximation.

### v2

@@ IV.2 Counting a set of specific transitions
$\beta$ $=C_{x_{A}\to y_{A}}W_{zz^{\prime}}\langle y_{A}|W_{AB}M|z\rangle\langle z^{\prime}|MW_{BA}|x_{A}\rangle,$ (59)
@@
$a_{ji}$ $=C_{x_{A}\to y_{A}}\sqrt{W_{z_{i}z_{i}^{\prime}}W_{z_{j}z_{j}^{\prime}}}\langle y_{A}|W_{AB}\tilde{M}|z_{i}\rangle\langle z_{j}^{\prime}|\tilde{M}W_{BA}|x_{A}\rangle,$ (62)
@@
Notice that these are all of dimension $r$. The rank of the matrices will be the number of transitions that are being monitored. This holds even if the state space itself is of much higher dimension. $\tilde{M}$ is defined as
@@
$P(\hat{Q}_{1},\ldots,\hat{Q}_{r})=\sum_{\rm all\penalty 10000\ perms}\mathrm{tr}\Big\{P_{1}a(P_{1}b)^{n_{1}}(P_{2}b)^{n_{2}}\ldots(P_{r}b)^{n_{r}}\Big\},$ (65)
@@
V Fluctuation theorem and thermodynamic uncertainty relation
The stochastic entropy production during an excursion was defined in Eq. (17). The probability of observing an entropy production $\sigma$ and the realization $q$ of an arbitrary anti-symmetric counting observable (such as a current) during an excursion starting from $x_{A}$ and ending in $y_{A}$ is given by Eq. (20):
$P_{x_{A}\to y_{A}}(q,\sigma)=C_{x_{A}\to y_{A}}\int\frac{d\xi_{q}d\xi_{\sigma}}{(2\pi)^{2}}$ $\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}W_{BA\bm{\xi}}|x_{A}\rangle$ (71)
$\times\penalty 10000\ e^{i(q\xi_{q}+\sigma\xi_{\sigma})},$
where $\bm{\xi}=\{\xi_{q},\xi_{\sigma}\}$ and
@@ V Fluctuation theorem and thermodynamic uncertainty relation
Due to the local detailed balance condition (16), and the symmetry imposed by time-antisymmetric weights of both $q$ and $\sigma$, the change of variable $\bm{\xi}\to\bm{\xi}^{\prime}\equiv\{-\xi_{q},-\xi_{\sigma}-i\}$ implies that the corresponding tilted matrix in Eq. (18) satisfies the special symmetry
$\mathbb{W}_{\bm{\xi}}^{\rm T}=\mathbb{W}_{\bm{\xi}^{\prime}}.$ (73)
Therefore, the probability with the new counting fields reads
$P_{x_{A}\to y_{A}}(q,\sigma)$ $=C_{x_{A}\to y_{A}}e^{\sigma}\int\frac{d\xi_{q}d\xi_{\sigma}}{(2\pi)^{2}}$ (74)
$\times\penalty 10000\ \langle x_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}W_{BA\bm{\xi}}|y_{A}\rangle e^{-i(q\xi_{q}+\sigma\xi_{\sigma})},$
where we have taken the transpose of the terms in bra-kets. The integral above is proportional to $P_{y_{A}\to x_{A}}(-q,-\sigma)$, except for the normalization constant, hence the fluctuation theorem
$\frac{P_{x_{A}\to y_{A}}(q,\sigma)}{P_{y_{A}\to x_{A}}(-q,-\sigma)}=e^{\sigma}\frac{C_{x_{A}\to y_{A}}}{C_{y_{A}\to x_{A}}}.$ (75)
The coefficient $C_{x_{A}\to y_{A}}/C_{y_{A}\to x_{A}}$ is the ratio between the probability of having an excursion $x_{A}\to y_{A}$ and having its reverse $y_{A}\to x_{A}$. This appears in Eq. (75) because the left-hand side denotes a conditional probability on both ends. If region $A$ has a single state, than the ratio $C_{x_{A}\to y_{A}}/C_{y_{A}\to x_{A}}$ becomes the one and we recover the typical “exchange” type fluctuation theorem. We numerically illustrate our result in Eq. (75) with simulated data from excursions in Fig. 2(a). We emphasize that this result holds for the realization $q$ of any counting observable as long as its weights are anti-symmetric (or equivalently the observable is anti-symmetric under time reversal).
Figure 2: Numerical illustration of the (a) excursion fluctuation theorem and (b) the thermodynamic uncertainty relation (TUR). For both plots we considered uniformly randomly generated stochastic transition matrices with entries $W_{ij}\in[0,2],\penalty 10000\ \forall\penalty 10000\ i\neq j$ and diagonal elements chosen such that $\sum_{i}W_{ij}=0$ to ensure normalization. Region AA was taken to have three states and region $B$ to also have three states. For (a) we fixed states $x_{A}=1$ and $y_{A}=3$ to be the initial and final states of the excursions. For (b) we consider an anti-symmetric counting observable randomly generated with entries $\nu_{ij}\in[0,2],\penalty 10000\ \forall\penalty 10000\ i>j$, with the constraint $\nu_{ij}=-\nu_{ji}$; and l.h.s and r.h.s represent respectively the left hand-side and right hand-side of Eq. (76).
Furthermore, it is known that a fluctuation theorem immediately implies a thermodynamic uncertainty relation Potts2019 ; Hasegawa2019 . In the case of our result in Eq. (75), we obtain
$\dfrac{{\rm var}(\hat{Q})_{x_{A}\to y_{A}}+{\rm var}(\hat{Q})_{y_{A}\to x_{A}}}{\big(E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}\big)^{2}}\geq\dfrac{1}{e^{[E(\hat{\Sigma})_{x_{A}\to y_{A}}+E(\hat{\Sigma})_{y_{A}\to x_{A}}]/2}-1},$ (76)
where averages and variances are conditioned over the initial and final excursion states. A proof of this result is provided in Appendix A.4, and numerical verification for the bound over randomly generated transition matrices is provided in Fig. 2(b). This uncertainty relation is valid at the excursion level, and it recovers the result of Ref. Hasegawa2019 when region $A$ has a single state $x_{A}=y_{A}$. A new insight follows directly from the excursion level fluctuation theorem. It reveals how fluctuations of a current along excursions $x_{A}\to y_{A}$ and their reverse $y_{A}\to x_{A}$ are bounded by entropy production. As with all thermodynamic uncertainty relations, it pinpoints a trade-off between precision and fluctuations. However, the result for excursions signifies in a concrete way the interplay between variance and averages of the forward and backward processes, which is not necessarily clear for the generic trajectory-level result.
@@ VI.1 Simplified absorption refrigerator
Quantum absorption refrigerators have been described in the literature for many years Mitchison2019 ; Levy2012 ; Levy2012a ; Correa2014 ; Correa2013 ; Cangemi2024 . Recently, they have received a surge of interest due to its experimental implementation aamirThermallyDrivenQuantum2025 ; blokQuantumThermodynamicsQuantum2025 ; Campbell2025 . The system consists of three qubits, each connected to their respective reservoirs which are assumed to be bosonic (at different temperatures), denoted $c$ (cold), $h$ (hot), and $w$ (work). The desired task is to cool down the cold qubit, so the temperatures are taken to obey $T_{c}<T_{h}<T_{w}$. Each qubit has a different energy gap given by $\omega_{\alpha}$ where $\alpha=c,\penalty 10000\ h,\penalty 10000\ w$. The three qubits interact resonantly (i.e. when $\omega_{c}+\penalty 10000\ \omega_{w}=\penalty 10000\ \omega_{h}$) through an effective three-body interaction, allowing the conversion of one excitation in the work and cold reservoirs into one in the hot reservoir, effectively cooling the cold reservoir in the process, see Fig. 3(a). The behavior of this system at the ensemble level is characterized by the cooling window, which provides bounds to the temperatures of the reservoirs such that the system works as a refrigerator:
@@
Figure 3: (a) Diagram of the three-qubit absorption refrigerator, each qubit is connected to its respective heat reservoir and has a different energy gap. Labels are $c$ for cold, $h$ for hot, and $w$ for work. The resonant interaction $g$ couples the states $|010\rangle\leftrightarrow|101\rangle$ and takes place only if $\omega_{c}+\omega_{w}=\omega_{h}$. (b) Energy landscape of the simplified three-qubit refrigerator model, where the arrows indicate the possible transitions. Colors indicate which reservoir induced the transition. Region AA consists of the ground state (gray) and $B$ of all excited states (colored). We adopt the convention “chwchw” to denote the states.
First, we qualitatively discuss a simplified version of the absorption refrigerator, where only five energy states are considered (in constrast with all eight states). In Ref. Fiusa2025 we discussed in detail results concerning three counting observables of interest: cold current, entropy production, and dynamical activity. In the next subsection, we promote a similar discussion for the full model. We denote the states by 0 (resp. $1$) representing that the qubit is in the ground state (resp. is excited), and we follow the convention chwchw. Region $A$ is taken to be spanned only by the state $|000\rangle$, where all qubits are in the ground state. $B$ is spanned by the other four states: $|001\rangle,\penalty 10000\ |010\rangle,\penalty 10000\ |100\rangle,\penalty 10000\ |101\rangle$ [Fig. 3(b)]. This is the order of states we adopt for our convention when denoting elements of the stochastic transition matrix. For example, the matrix element $W_{14}$ represents the transition rate $|100\rangle\to|000\rangle$. In the parameter regime of Ref. aamirThermallyDrivenQuantum2025 , the classical master equation (1) approximates the quantum dynamics very well provided that the coupling $g$ is replaced by some classical $g^{\prime}$, see Appendix B.
The transition among states come from two sources, the reservoirs whose transitions rates are described by $\Gamma_{\alpha}n_{\alpha}$ for injections and $\Gamma_{\alpha}(n_{\alpha}+1)$ for extractions, and the three body interaction $g^{\prime}$ that couples the states $|010\rangle\leftrightarrow|101\rangle$ provided that $\omega_{h}=\omega_{c}+\omega_{w}$. Since the transitions are induced by a single reservoir, we omit the index $\ell$ in our discussion. $n_{\alpha}=\left[\exp({\omega_{\alpha}}/T_{\alpha})-1\right]^{-1}$ denotes the occupation number and $\Gamma_{\alpha}$ is the coupling strength between the qubit and the reservoir. The stochastic transition matrix (without the diagonal elements) is then given by
@@
Given the topology of the states and transitions in $B$, this model inherits a nice feature when it comes to constructing the counting observables from Eq. (15). Namely, the thermodynamical currents such as cold, hot and work currents can only have three possible values per excursion. To see this, consider the cold current that is constructed by setting the weights $\nu_{xy}=\pm 1$ in Eq. (15) if the transition $x\rightarrow y$ removes (resp. adds) an excitation from (resp. into) the cold reservoir. The three possibilities for the cold current are: a success where cooling took place and the cold current is $\hat{J}_{c}=1$, e.g. $|000\rangle\to|100\rangle\to|101\rangle\to|010\rangle\to|000\rangle$; a disaster where heating took place the cold current is $\hat{J}_{c}=-1$, e.g. $|000\rangle\to|010\rangle\to|101\rangle\to|001\rangle\to|000\rangle$, and a bounce/fail where $\hat{J}_{c}=0$, e.g. $|000\rangle\to|100\rangle\to|000\rangle$. This feature also arises for the entropy production, where the positive and negative contributions are related via fluctuation theorems (75).
@@ VI.2 Full absorption refrigerator
Now we turn our attention to the full model, with all eight states: $|000\rangle,\penalty 10000\ |001\rangle,\penalty 10000\ |010\rangle,\penalty 10000\ |011\rangle,\penalty 10000\ |100\rangle,\penalty 10000\ |101\rangle,\penalty 10000\ |110\rangle,\penalty 10000\ |111\rangle$. The three reservoirs are bosonic with occupation number $n_{\alpha}$ and each qubit is coupled to its reservoir with coupling strength $\Gamma_{\alpha}$, where $\alpha=c,\penalty 10000\ h,\penalty 10000\ w$. All elements of the transition matrix therefore have the form $\Gamma_{\alpha}n_{\alpha}$ and $\Gamma_{\alpha}(n_{\alpha}+1)$, except for the transition $|101\rangle\leftrightarrow|010\rangle$, which is denoted $g^{\prime}$. The stochastic transition matrix (without the diagonal elements) is then given by
@@
First, we begin with the cold current which is constructed by setting the weights $\nu_{xy}=\pm 1$ in Eq. (15) if the transition $x\rightarrow y$ removes (resp. adds) an excitation from (resp. into) the cold reservoir. To study the cold current we take region $A$ to be spanned by the states $|000\rangle,\penalty 10000\ |010\rangle,\penalty 10000\ |001\rangle,\penalty 10000\ |011\rangle$. This is the natural choice for the full refrigerator model, because now the excursions start whenever an excitation enters the cold qubit and end whenever the excitation leaves. This also assures the nice property that the cold current may only assume three possible values per excursion. Indeed, following the counting field expansion [see Eq. (20)], the probability of the values for the cold current is written as
$P(\hat{J}_{c})$ $=\int_{-\infty}^{\infty}\dfrac{d\xi}{2\pi}\left(P_{\textrm{suc}}\penalty 10000\ e^{i\xi}+P_{\textrm{dis}}\penalty 10000\ e^{-i\xi}+P_{\textrm{fail}}\right)$ (80)
$=P_{\rm suc}\delta(\hat{J}_{c}-1)+P_{\rm fail}\delta(\hat{J}_{c}-0)+P_{\rm dis}\delta(\hat{J}_{c}+1),$
@@
since the span of the cold current is discrete, the integration results in the Fourier transform of the Dirac deltas centered around each possible value of $\hat{J}_{c}$. Note that the coefficients in the expansion of the counting field are the probabilities that the counting field variable has picked up a particular phase. In other words, $P_{\textrm{suc}}$ is the probability that the cold current has value $\hat{J}_{c}=1$ in an excursion. Likewise, $P_{\textrm{dis}}$ is the probability that the cold current has value $\hat{J}_{c}=-1$ and $P_{\textrm{fail}}$ the probability of $\hat{J}_{c}=0$. Fig. 4(a) shows the behavior of the probabilities as a function of the cold bath occupation number $n_{c}$, where it is clear that cooling failures (bounces) dominate the dynamics. The value where $P_{\rm suc}$ becomes larger than $P_{\rm dis}$ is precisely the value of $n_{c}$ where the cooling window begins. In Fig. 4(b), we show the covariance between the cold current $\hat{J}_{c}$ and the excursion duration $\hat{T}$ as a function of the cold bath occupation $n_{c}$. Interestingly, longer excursions tend to have larger $\hat{J}_{c}$ in the cooling window, which is not necessarily the case when $\langle\hat{J}_{c}\rangle>0$, and once again, a vanishing covariance pinpoints the value of $n_{c}$ where the cooling window begins.
Equation (80) illustrates another feat of the excursion framework. This is a new paradigm for analyzing the performance of cooling/heating protocols beyond the steady state. The steady state cold current can be written as function of the probabilites as follows
$J_{c}^{\rm ss}=\dfrac{P_{\rm suc}-P_{\rm dis}}{\mu}.$ (81)
@@
This result can also be recovered by computing a generic classical current by solving for the steady state in Eq. (2). The steady state current takes into account only the difference between the probability of successes and disasters, overlooking how dominant failures are. This raises questions for designing useful thermodynamic protocols, e.g. a protocol may have a large cooling capacity $P_{\textrm{suc}}\gg P_{\textrm{dis}}$ but have low cooling precision $P_{\textrm{fail}}\gg P_{\textrm{suc}}$. Non-trivial tradeoffs emerge from this analysis.
@@
For the next two counting observables, we go back to the case where region $A$ is spanned only by the state $|000\rangle$. This allow us to use the results from Eqs. (46) and (47) to characterize fluctuations. Entropy production $\hat{\Sigma}$ was constructed in Eq. (17), and we highlighted its importance in our framework when we discussed fluctuation theorems in Sec. V. Likewise, the excursion activity $\hat{\mathcal{A}}$ was constructed in Eq. (48) and we have provided analytical formulas for its expectation value in Eq. (55) and variance in Eq. (IV.1). In Figs. 5(a),(b) we characterize the diffusion coefficient (47) in terms of its components, both for the entropy production $D_{\Sigma}$ and the dynamical activity $D_{\mathcal{A}}$. In the former, the fluctuations are strongly dominated by the variance term of $\hat{Q}=\hat{\Sigma}$; whereas in the latter there is a clear competition between different contributions, with $E(\hat{\mathcal{A}})^{2}$ being the most relevant term. This result sheds light on the very nature of noise in FCS, where the dominant contribution can change depending on the counting observable. $\hat{\Sigma}$ is an observable that can have positive and negative values related by fluctuation theorems, whereas $\hat{\mathcal{A}}$ is strictly positive. The stark contrast of behavior on the two observables shows the versatility of the excursion framework to characterize fluctuations.
@@
Figure 5: Diffusion coefficient and its decomposition for (a) entropy production, and (b) dynamical activity as a function of $g$, with $n_{c}=0.1$ and $\Gamma_{c}=0.1$. For all plots, we fixed $\Gamma_{h}=50,\ \Gamma_{w}=0.5,\ n_{h}=0.005,\ n_{w}=0.5$. Parameters are well in accordance with the parameter regime of aamirThermallyDrivenQuantum2025 and the classical approximation.